OpenStudy (anonymous):

a ladder is leaning against a building so that the distance from the ground to the top of the ladder is 2feet less than the length of the ladder.Find the length of the ladder if the distance from the bottom of the ladder to the building is 10feet

OpenStudy (wolf1728):

L² = (L-2)² +10² L² = L² -4L + 4 + 100 4L = 104 L = 26 Feet